Express the confidence interval (67.8%,82.4%)(67.8%,82.4%) in the form of ˆp±MEp^±ME.

______% ±± _______%
Answer #1

Solution :

Point estimate = \hat p = (Lower confidence interval + Upper confidence interval ) / 2

Point estimate = \hat p = (0.678+0.824) / 2=0.751

Margin of error = ME = Upper confidence interval - \hat p = 0.824-0.751=0.073

In the form of :

\hat p ±ME

0.751 ±0.073